When 27.4 mL of 0.500 M H2SO4 is added to 27.4 mL of 1.00 M KOH in a coffee-cup calorimeter at 23.50°C, the temperature rises to 30.17°C. Calculate ΔH of this reaction. (Assume that the total volume is the sum of the individual volumes and that the density and specific heat capacity of the solution are the same as for pure water.) (d for water = 1.00 g/mL; c for water = 4.184 J/g·°C.) kJ/mol H2O
Answer #1

temprature changes from 23.50c to 30.170c therefore \DeltaT = 30.17 - 23.5 = 6.67

total volume of soln = volume of H2SO4 + volume of KOH =27.4+27.4=54.8 ml

density of solution is 1gram per ml therefore 54.8ml solution = 54.8 gram solution =mass of solution=54.8gram

specific heat of solution =specific heat of pure water = 4.184 J/g

we know formula \DeltaH = mass of solution \times specific heat \times\DeltaT

substitute values in above equation

\DeltaH = 54.8 \times 4.184 \times 6.67 = 1529.3 J

\DeltaH for reactiion = 1529.3 J